OpenStudy (anonymous):

H12P1 section c plzzz...

OpenStudy (anonymous):

c) Q: v=5v A=100, load resistance =500 Solve two equations to find value of vx vx/R=(vOut–vx)/500,with R=answer from Part1. vx=5–vOut/100 Solving two equations we get Vx Current = Vx / R

OpenStudy (anonymous):

I=Vout/R1+R2 R1=R2